• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M Friend

    Log entry time 10:29:59 on October19,2009

    Entry number 295672

    keyword=Compton instabilities

    Compton rates and the current in the Compton photon detector PMT base have been unstable since 19:30 last night. The green line is the current in the base, and if it were stable it would be completely flat. Here it is not flat. It was unstable all night.

    FIGURE 1